OpenStudy (anonymous):

given sin (a)=(-43/45) and tan (a) > 0, determine the value of cos (a)

OpenStudy (y2o2):

sin(a) is -ve and tan(a) is +ve so (a) is in the 3rd Quad. \[\sin^{-1}({43 \over 45}) = 72.85379\] a = 180+72.85379 = 252.85379

OpenStudy (y2o2):

you can also use the identity of \[\sin^2 + \cos^2 =1\] it will be a lot better

OpenStudy (anonymous):

so whts the value of cos a?

OpenStudy (y2o2):

\[\cos(a) = -{4\sqrt{11} \over 45}\]

OpenStudy (anonymous):